OpenStudy (anonymous):

find the integral of [(x^2)divided by squareroot of (x+2)] dx

OpenStudy (anonymous):

\[\int \frac{x^2}{\sqrt{x+2}}dx\]

OpenStudy (anonymous):

is that Right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmm you can add 4 and negative 4 in numerator hmm let me do this

OpenStudy (anonymous):

\[ \frac{x^2 +4 -4 }{\sqrt{x+2}} \implies \frac{(x+2)(x-2)}{\sqrt{x+2}} + \frac{4}{\sqrt{x+2}}\]

OpenStudy (anonymous):

sorry no medal.. I need the way on how to answer this

OpenStudy (anonymous):

oh don't worry about that Lol

OpenStudy (anonymous):

hmm I have a better way now put \( x+2 = t^2\)

OpenStudy (anonymous):

\[x = t^2 - 2 \] \[dx = 2x* dt \] \[\frac{x^2}{\sqrt{t^2}}\implies \frac{(t^2-2)^2}{t}\]

OpenStudy (anonymous):

Now you can do it

OpenStudy (anonymous):

Correction :\[dx = 2t* dt\]

OpenStudy (anonymous):

\[2\int \frac{\cancel{t}(t^4 + 4 - 4t) }{\cancel t}dt\]

OpenStudy (anonymous):

\[2 \int( t^4 +4 -4t^2)dt\]